PAE activity log

This page concerns the logging of the PAE activities. The instrument configuration was frozen on March 31st, 2017, and all subsequent activities have to be recorded, through these pages.

To record a new activity, please fill up this form

Hereafter are all the logged activities, sorted by date

Date
Name
Activity
Affected Subsystems
Details
2017-04-20
Roberto Cirami
Test
System
A whole set of tests to check all the instrument subsystems have been executed in preparation of the software PAE.
2017-04-19
Roberto Cirami
Test
System
A whole set of tests for checking all the instrument subsystems have been executed in preparation of the software PAE.
2017-04-19
Roberto Cirami
Configuration change
ICS
The shutter instrument workstation software has been updated in order to allow to reset a PLC error condition if the NGC try to take an exposure with no shutter selected.
2017-04-18
Giorgio Calderone
Test
System
I started the procedure to run a thorough test of all available ESPRESSO systems.
This operation failed last time it has been executed since the FS3 camera was not reachable.

It will take approximately 3-4 hours to run all the tests. I will describe the test outcomes in the next log entry.

The tests can be interrupted at any time, just drop me an email....
2017-04-18
Giorgio Calderone
Test
Front End
The FS3 camera (IP address: 192.168.66.155) do not responds to ping queries, as if it has been switched off, or is not connected to the network
2017-04-14
Roberto Cirami
Configuration change
ICS
The ESPRESSO software has been updated to the latest version (revision 296414). Before doing the update the DETDATA directory has been copied in /diskb/Backup_DETDATA/2017-04-14. Some tests have been performed that involved ICS, scientific detector and TCCDs.
2017-04-14
Pepe Francesco
Operation
Cryo-vacuum system
Thursday 2017-04-13
- Alarm on 'bad vacuum' received by Ian Hughes
- F. Pepe checked and 'reset' the alarm on the system at OG around 10 pm.
- Cryo-Pump 1 was set on 'normal' mode (was on engineering mode before)
- A few hours later, the emergency pumping started automatically

Friday 2017-04-14
- From remote (VNC) F. Pepe stopped the emergency pumping, since Cryo-Pump 1 must not be operational when Cryo-Pump 2 is.

Remark:
- VV still has a leak on the blue flange, as reported earlier. The leak is of the order of 2E-4 per day. The leak shall be repaired at a later occasion, since not critical.

2017-04-10
Francesco Pepe
Test
System , Science SW
Starting on Friday 7 of April afternoon to Monday 10 of April we have performed a long sequence of TAL1-TAL2 and DARK exposures. The goal was to:

- Test the system operability on a long time series
- Perform repeated DARK exposures
- Test the 'drift-measurement' recipe
- Test the absolute and relative (fiber-to-fiber) stability of the spectrograph

The test was conducted successfully and the data are being processed.
2017-04-13
ian hughes
Test
Spectrograph, Blue Camera, Thermal system, Cryo-vacuum system
We performed a blackout test on the instrument. The system was in operation mode before cutting the power and the pumping was enable. The power cut was simulated by unplugging the main power cable on the wall. this shut down everything except for the computers. Following the procedure, we restart every cabinets and devices and 15 minutes later, the system was recovered. in the afternoon, a leak detection was performed on the vacuum vessel in order to check if the leak was repaired. unfortunately the leak is still there, but slightly less than before. The actual leak value is 4x10-8 mbar*l/s. The system has been put back in automatic mode for the Easter w-e.
2017-04-12
Francesco Pepe
Operation
System , Cryo-vacuum system
At 7 pm I have:

- Stopped VV pumping (engineering mode)
- Verified that Cryo-Pump 2 was at nominal cooling temperature (110K) and operational (valve on VV open)
- Started cooling of Cry-Pump 2

The pressure of the VV after closing the pumping VV was 9E-6 mbar at 19.00 hour.
2017-04-12
Francesco Pepe
Configuration change
Cryo-vacuum system
The screws of the VV vacuum flange on which the leak has been detected earlier on were found to be poorly tighten. Due to the nature of this flange, which is mounted from inside the vacuum vessel, the leak might be cause by this fact. After tighten the screws the pressure raise of the WARM VV was measured to about 4E-5 mbar per hour, i.e. less than 1E-3mbar per day. This is both a lower value compared to the raising rate measured previous while using the cryostat-pumps and a value three time lower than the HARPS vessel outgassing. We think therefore that the vacuum leak has been fixed. After re-generation, we (Nicolas Haddad, Ian Hughes, Francesco Pepe) will start the cryostat-pumping (#2) and see whether the VV pressure remains stable.
2017-04-12
Francesco Pepe
Configuration change
Cryo-vacuum system
As reported by Ian Hughes earlier, a leak was found on the blue detector cryostat and fixed. The confirmation for having fixed the leak was obtained today by stopping the pumping and observing how the pressure in the blue dewar evolved. When stopping the pump and closing the valve at about 10 am, the pressure was about 1.6E-6 mbar. It raised somewhat but at 5 pm it seemed quite stable around 2.43E-6 mbar.

The pressure seems does performing well despite the fact that no re-generation of the sorption pump has been performed after the leak repair.
2017-04-11
Ian Hughes
Maintenance
Blue Camera, Thermal system, Cryo-vacuum system
A more accurate leak detection has been performed on the vacuum vessel Blue flange and the leak has ben located. It seems to be between the big flange and the intermediate flange of the blue cryostat. An other leak detection has been performed on the Blue Detector and it appears that a over-pressure valve was leaking. This leak has been solved by rotating the top part of the valve which is spring-loaded sealed. the systems needs to recover now and tomorrow a new vacuum test will be preformed.
2017-04-10
Ian Hughes
Maintenance
Spectrograph, Blue Camera, Cryo-vacuum system
Due to a abnormal raise of the pressure inside the vacuum vessel when connected to the cryo-pump, we performed a leak detection with Nicolas Haddad and with the help of Jean-Louis. We detected a leak around the flange of the Blue cryostat. Tomorrow, this will be investigated more deeply. Jean-Louis will be here.
2017-04-07
Ian Hughes
Operation
Spectrograph, Blue Camera, Thermal system, Cryo-vacuum system
Replaced the empty tank of LN2 by a full one. Has to be done twice a week.
2017-04-06
Francesco Pepe
Test
System
We have executed a full set of calibrations. The LDLS flat-field lamp was used with the colour-balancing filter. At the same time, I have determined the default filter setting and exposure times for a correct flux level.

Calibration set:
cal_order: TEXP = 5 s, NDA = 2.0, NDB = 1.7, achieved max flux 55'000 ADU
cal_FF: TEXP = 5 s, NDA = 2.0, NDB = 1.7, achieved max flux 55'000 ADU
cal_wave with TAL1-TAL2: TEXP = 40 s, NDA = 0, NDB = 0
cal_wave with FP-FP: TEXP = 40 s, NDA = 2, NDB = 1.7
cal_wave with TAL1-FP: TEXP = 40 s, NDA = 0, NDB = 1.7
2017-04-03
Francesco Pepe
Configuration change
Calibration System
In order to lower the dynamical range between the red and blue end of the blue-camera spectrum. I have inserted a filter box and a 200 µm fiber between the LDLS fiber and the calibration-unit fiber of the FF lamp. The filter box was populated with a thermal filter cutting above 780 nm and an interference filter by TFP reproducing the inverse of a BB-curve with a transmission decreasing by a factor of 10 to 15 between 400 nm and 800 nm. After insertion of these filters the dinamic range was reduced from 10 to 4 and is well belanced in color.
The configuration change is provisional in the sense that
- The filter must be validated also by the red camera.
- The documentation must be updated if the filter is kept as final solution.

It must be noted that calibration passed QC also without the filter, but exposure time tuning is more critical to avoid saturation in the red and have enough flux in the blue.

Comments are closed.